# Leveraging Photometry for Deconfusion of Directly Imaged Multi-Planet Systems

### v1

@@ 2.3.3 Confused System Orbit Options
The number of confused orbits per simulated system are listed in the three right columns of Table 2. The system that we will use as an example to illustrate the updated ranking process, system ten, has two confused orbit options for the low inclination case, two confused orbit options for the medium inclination case, and two confused orbit options for the high inclination case. The orbit options for each inclination grouping for system ten are shown in Figure 4.
For each set of confused orbit options (i.e., for each system), we apply the new photometry likelihood ranking scheme by simulating the expected brightness of each detection given the orbital parameters fit by the deconfuser. The likelihoods for each confused system are compared to rank the matched orbits in order of most to least likely for each system. Rankings for the low, medium, and high inclination cases of system ten are given in Table 3. Table 3 also provides the semimajor axis, eccentricity, and inclination, and rms error of each planet’s orbit option as fit by the deconfuser.
@@
Figure 4: Confused orbit options for the low (a and b), medium (c and d), and high inclination (e and f) detections for system ten (detections shown in Figure 3). The rankings for each of the confused options across the inclination groupings are given in Table 3. The orbit fits from each possible set of options are compared to the true simulated system’s orbits in Figure 5 for each inclination group.
Table 3: System ten photometry ranking results. Orbital parameters and updated rankings for the confused orbit options returned by the deconfuser for the low, medium, and high inclination system ten. The confused orbit options are shown in Figure 4. RMS errors for each individual orbit fit by the deconfuser without photometry are also listed. Results for all simulated systems are provided in a supplemental file.
“Confused” orbit parameters
Incl. System Planet no. $a$ $e$ $i$ Deconfuser RMS Orbit System System
Group option (AU) (°) fit error likelihood likelihood ranking
@@ 3 Results
Here we present findings regarding the performance of the updated ranking scheme using planetary phase and brightness. The results are based on a subset of highly confused simulated systems, while assuming that all planets are visible at all times. Simulated detections are passed through the deconfuser and photometry model without knowledge of which detection corresponds to which planet. For each of the simulated systems that we apply the photometry ranking to, the deconfuser finds orbit solutions which fit the detections, but is unable to identify a single solution for each system given only their positional information (i.e., confusion). We use test case system ten as an illustrative example in this section. Photometry ranking results for all systems and inclinations are provided in a supplemental file.
In Section 3.1 we describe how the updated photometry ranking performs for the subset of highly confused systems as a function of inclination. In Section 3.2 we validate the orbits fit by the deconfuser through a comparison with fits from orbitize!.
@@ 3.1 Photometry Ranking for Deconfusion
Here we describe the improvement in deconfusion for the example systems presented that is brought by including photometry with our new ranking scheme. In the case of system ten, photometry aids in deconfusing the system by providing an updated ranking to each set of orbits for the system. We compare the top-ranked orbit match to the true system by calculating the percent difference between individual orbit parameters for each planet. The orbit fits which more closely resemble the true simulated system parameters for system ten were ranked highest across all inclination groupings. “Closely” here refers to the minimum percent difference between orbit parameters across all confused orbit matches.
For the low inclination case (detections in Figure 3, panel A), the deconfuser originally returned two equally likely fits for the system. The first option (option A) depicted a relatively co-planar system with low eccentricities for all planets and inclinations of less than 28 (panel a in Figure 4). Option B included a similar fit for planet one in the system, but planets two and three were fit with high inclination orbits and eccentricities of 0.1 (panel b in Figure 4). The photometry ranking scheme determined that option B was statistically less likely based on the detected planetary brightnesses. The same is true for the medium (Figure 4, Panels c and d) and high inclination (Figure 4, Panels e and f) cases of system ten. The top ranked system options are shown compared to the true simulated system orbits in Figure 5.
Figure 5: Top ranked orbit options (dashed lines) for system ten compared to the true simulated system’s orbits (solid black lines) for each system: (a) low inclination, (b) medium inclination, and (c) high inclination.
For all simulated systems in the low inclination grouping, the photometry ranking scheme ranked the “best” confused option (in terms of semimajor axis) higher in all but 30% of the ten cases. “Best” here refers to the confused system match that has the minimum percent difference between each planets’ orbital parameters; “worst” refers to a match that does not have the minimum percent differences out of all confused options. For the medium and high inclination groupings, the photometry ranking scheme also ranked the best system option higher in all but 40% of cases. Figure 6 presents the likelihood results for all test cases examined in this work, with different marker shapes corresponding to the photometry ranking assigned to each orbit fit in the system. We emphasize that prior to ranking with photometry, all of the systems tested here were confused and these are improvements with a 30% to 40% error rate. With only relative astrometry, the deconfuser was unable to choose just one set of orbits within the astrometry error tolerance. We present a discussion of these results in Section 4 and directions for further improving deconfusion with photometry in Section 5.
@@
3.2 Validating deconfuser Orbit Fits
[…]
To compare the simulated system detections generated in this work, we first convert the on-sky positions from Cartesian coordinates to relative right ascension and declination. We follow the conversion process employed by orbitize!, which was initially presented in Green (1985) and is shown in Equations 9 and 10.
$\Delta\mathrm{R.A.}=\pi a(1-e\cos E)\Big[\cos^{2}{\tfrac{i}{2}}\,\sin(f+\omega_{p}+\Omega)\\ -\sin^{2}{\tfrac{i}{2}}\,\sin(f+\omega_{p}-\Omega)\Big]$ (9)
$\Delta\mathrm{decl.}=\pi a(1-e\cos E)\Big[\cos^{2}{\tfrac{i}{2}}\,\cos(f+\omega_{p}+\Omega)\\ +\sin^{2}{\tfrac{i}{2}}\,\cos(f+\omega_{p}-\Omega)\Big]$ (10)
The orbital parameters are defined in Table 4.
MCMC and other Bayesian processes are undefined if the error is zero. Therefore, we must include some error on the relative astrometry in order to employ MCMC for orbit-fitting. We apply a small random Gaussian error to the astrometry with $\sigma=$ 1 mas. This is comparable to current real-life astrometry error (e.g., Wang et al. 2016).

### v2

@@ 2.3.3 Confused System Orbit Options
The number of confused orbits per simulated system are listed in the three right columns of Table 2. The system that we will use as an example to illustrate the proof-of-concept, system ten, has two confused orbit options for the low inclination case, two confused orbit options for the medium inclination case, and two confused orbit options for the high inclination case. The orbit options for each inclination grouping for system ten are shown in Fig. 4.
For each set of confused orbit options (i.e., for each system), we apply the new photometry likelihood ranking scheme by simulating the expected brightness of each detection given the orbital parameters fit by the deconfuser. The likelihoods for each confused system are compared to rank the matched orbits in order of most to least likely for each system. Rankings for the low, medium, and high inclination cases of system ten are given in Table 3. Table 3 also provides the semimajor axis, eccentricity, inclination, and rms error of each planet’s orbit option as fit by the deconfuser.
@@
Figure 4: Confused orbit options for the low (a and b), medium (c and d), and high inclination (e and f) detections for system ten (detections shown in Fig. 3). The rankings for each of the confused options across the inclination groupings are given in Table 3. The orbit fits from each possible set of options are compared to the true simulated system’s orbits in Fig. 5 for each inclination group.
Table 3: System ten photometry ranking results. Orbital parameters and updated rankings for the confused orbit options returned by the deconfuser for the low, medium, and high inclination system ten. The confused orbit options are shown in Fig. 4. RMS errors for each individual orbit fit by the deconfuser without photometry are also listed. Results for all simulated systems are provided in a supplemental file.
“Confused” orbit parameters Deconfuser
Incl. System Planet $a$ $e$ $i$ RMS Orbit System System
group option no. (AU) (°) fit error likelihood likelihood ranking
@@ 3 Results
Here we present findings regarding the performance of the updated ranking scheme using planetary phase and brightness. The results are based on a representative set of corner case systems systems, while assuming that all planets are visible at all times. Simulated detections are passed through the deconfuser and photometry model without knowledge of which detection corresponds to which planet. For each of the systems that we apply the photometry ranking to, the deconfuser finds orbit solutions which fit the detections, but is unable to identify a single solution for each system given only their positional information (i.e., confusion). We use test case system ten as an illustrative example in this section. Photometry ranking results for all systems and inclinations are provided in a supplemental file.
@@ 3.1 Photometry Ranking for Deconfusion
Here we describe the improvement in deconfusion for the representative systems presented that is brought by including photometry with the new ranking scheme. In the case of system ten, photometry aids in deconfusing the system by providing an updated ranking to each set of orbits for the system. We compare the top-ranked orbit match to the true system by calculating the percent difference between individual orbit parameters for each planet. The orbit fits which more closely resemble the true simulated system parameters for system ten were ranked highest across all inclination groupings. “Closely” here refers to the minimum percent difference between orbit parameters across all confused orbit matches.
For the low inclination case (detections in Fig. 3, panel A), the deconfuser originally returned two equally likely fits for the system. The first option (option A) depicted a relatively co-planar system with low eccentricities for all planets and inclinations of less than 28 (panel a in Fig. 4). Option B included a similar fit for planet one in the system, but planets two and three were fit with high inclination orbits and eccentricities of 0.1 (panel b in Fig. 4). The photometry ranking scheme determined that option B was statistically less likely based on the detected planetary brightnesses. The same is true for the medium (Fig. 4, Panels c and d) and high inclination (Fig. 4, Panels e and f) cases of system ten. The top ranked system options are shown compared to the true simulated system orbits in Fig. 5.
Figure 5: Top ranked orbit options (dashed lines) for system ten compared to the true simulated system’s orbits (solid black lines) for each system: (left) low inclination, (center) medium inclination, and (right) high inclination.
For all test case systems in the low inclination grouping, the photometry ranking scheme ranked the “best” confused option (in terms of semimajor axis) higher in all but three of the ten cases. “Best” here refers to the confused system match that has the minimum percent difference between each planets’ orbital parameters; “worst” refers to a match that does not have the minimum percent differences out of all confused options. For the medium and high inclination groupings, the photometry ranking scheme also ranked the best system option higher in all but four of the ten cases. Fig. 6 presents the likelihood results for all test cases examined in this work, with different marker shapes corresponding to the photometry ranking assigned to each orbit fit in the system. We emphasize that prior to ranking with photometry, all of the systems tested here were confused and these are improvements for more than half of the representative highly confused cases. With only relative astrometry, the deconfuser was unable to choose just one set of orbits within the astrometry error tolerance. We present a discussion of these results in Sec. 4 and directions for further improving deconfusion with photometry in Sec. 5.
@@
4 Discussion
Here we present a more detailed discussion of the results, including how well the orbits are fit compared to the truth and the implications of the results for direct imaging missions. Our proof-of-concept suggests that including photometry benefits deconfusion for the example systems in this work, across all inclination ranges. Across the low, medium, and high inclination cases simulated here, the photometry likelihood ranking scheme enabled the separation of confused orbit options for each of the systems. Photometry aided in choosing the better orbit fit out of the confused options in more than half of the cases across inclination groups.
4.1 Photometry for Deconfusion
There are several contributing factors which influence the utility of photometry for deconfusing orbit options in these example cases. The first is due to the statistical nature of photon detections when observing exoplanets and the limited number of photons which will be received when imaging at high contrast levels. The likelihoods that are used as the ranking metric in this work are derived from a distribution of potential electron counts at the detector for each planet detection, which is a combination of a Poisson distribution, Gamma distribution, and Gaussian distribution (see detailed description in Sec. 2.2). The combination of statistical distributions and low number of photons per observation may result in variations of estimated count rate per detection, which influences the likelihood value. This explains the potential for variation in likelihoods on a per-orbit level. That is, if there are two system options for which one planet was fit with the same orbit in both, the options will return different likelihood values between the two confused systems. This may ultimately affect the resulting system ranking for better or worse, even though all likelihood results fall within the distribution of possible expected values.
Another commonality between systems for which the “worst” option is chosen appears when planet detections are close to each other in the plane of the sky. In system one, the detections of the two inner planets appear to be less than 0.1 AU from one another in the plane of the sky at each inclination. Systems four and five also fall into this category. In system four, the second detections of the inner two planets appear closer than 0.3 AU (or 0.03′′) of one another. For this reason, the difference in likelihoods between the low inclination cases are relatively small as the predicted phase angles are similar (in combination with the limited number of received photons) and spatially close together. Figure 6 shows the likelihoods for all systems and ranked options, and is shaded based on the number of planets per ranking for which the semimajor axis was the “best” fit for the system (i.e., minimum percent difference between true and fit semimajor axes). The difficulty in separating options with detections with small separations can be seen in Fig. 6 for systems one, four, and five, where the differences between system likelihoods are small. Section 5 expands on methods that could investigate reducing the confusion caused by closely separated detections.
Figure 6 introduces some key points regarding photometry for orbit discrimination. For the example systems presented here, it is clear that photometry shows promise for supporting orbit differentiation better than geometric orbit-fitting alone. Previously, each of the system options presented were closely ranked using only relative astrometric information. Likelihoods assigned to these systems by the photometry ranking scheme allow for a refined ranking of orbit combinations from the deconfuser. The photometry ranking scheme generally provides more distinction between orbits for the high inclination systems in this work, compared to lower inclination, because there is more phase variation in these observations. This can be seen by the higher scatter in likelihood values in the high inclination cases for the systems presented here (e.g., systems four, five, and nine).
4.2 The Role of Geometric Albedo
We briefly comment here on the role of geometric albedo for modeling photometry, as this is currently a constant parameter in our simulations. The choice in geometric albedo ($A_{g}$) for modeling the photometry will affect the overall likelihood values if the albedo is guessed incorrectly. A model albedo that is larger than the observed albedo would require a lower phase angle, decreased planet-star separation, or larger planet radius to produce the observed signal; a model albedo that is smaller than the observed albedo would require the opposite. An incorrectly guessed albedo might preference a fit with larger/smaller planet-star separation if the model albedo is less/greater than the truth, if such an orbit-fit has been returned by the deconfuser. However, if all other parameters are held constant, the general trend of a planet’s phase variation remains the same for all values of $A_{g}$. Because the photometry ranking scheme is choosing between and comparing orbits that have already been fit to detections, rather than influencing the orbit fitting process, we do not expect an incorrect guess for the geometric albedo to drastically change the performance of the photometry ranking.
4.3 Phase Function and Albedo Assumptions
A limitation of this work lies in the choice of a Lambertian phase function and the constant albedo to model planetary reflectance. The assumption of Lambertian reflectance approximates the shape of a planetary phase function relatively well for the type of planets and temperature regimes that would be probed by a reflected light direct imaging mission. Strong asymmetries in the reflected light curves of cool, mature planets are less likely because they are receiving relatively uniform irradiation from their host star (i.e., they are not tidally locked like known short-period transiting planets; e.g., Refs. Zellem et al. [71], Parmentier et al. [51]). However, Solar System phase curves have been shown to deviate from Lambertian assumptions (e.g., Refs. Mayorga et al. [45], Mallama and Hilton [37], Mayorga et al. [44]).
Observations of Solar System planets have also demonstrated that real planetary albedos vary with wavelength and time, driven by factors such as clouds, aerosol scattering, surface features, and seasons. Reference Mallama and Hilton [37] discusses some of these results in detail. For example, analysis of models and observations show that the Earth may exhibit magnitude variation ranging from 0.12 to 0.76 between cloud-free and cloudy cases (e.g., Refs. Tinetti et al. [66], Stephens et al. [63], Mallama and Hilton [37]). The phase curve of Venus shows a reversal in brightness at high phase angles, which has been attributed to forward scattering of sulfuric acid droplets in the atmosphere [39]. The giant planets in the Solar System exhibit albedo variation on a smaller scale at visible wavelengths, compared to the terrestrial planets [37]. Jupiter observations near full phase show magnitude variations at the level of a few hundredths due to changes in the cloud bands [38].
These complexities are not yet included in this work, but we describe future directions in Sec. 5 to incorporate more realistic reflectance behavior. In the scheme of the likelihood ranking, the true solution will always lie in the sample of possible solutions that are calculated, if the forward model assumptions are correct. However, one might reject the best-fit orbit solution if the model lacks fidelity. For example, in the case where the true planetary reflectance is non-Lambertian.
4.4 Exo-moons and Exo-rings
Other factors that may influence exoplanetary photometry are moons and rings, which could have a negative impact on the photometry likelihood ranking scheme. Rings alter the spectrum of a planet and would introduce asymmetries to the phase curve in reflected light [1, 15, 10]. Previous work has suggested that an icy ring around an Earth-like planet could result in a brightness 100 times that of a ringless planet [1]. In such a case, additional observations at longer wavelengths would be important for breaking the albedo-radius degeneracy and inferring the presence of rings.
For exomoons, previous work indicates that the influence on the spectrum of an Earth-like planet is minimal near visible wavelengths [33]. Their study concluded that detecting a moon-Earth mutual event at 550 nm with a bandwidth of 20% with an 8-meter class telescope would require around two to twenty mutual events. Such cases where observations measure anomalous photometry due to some influencing factors would need to be considered on a case-by-case basis. Using a more complex phase function and incorporating additional filters or polarimetry would likely be more beneficial for ranking such systems where phase curve asymmetries are being induced by material in the circumplanetary environment.
4.5 Effect of Planetary Motion on Astrometry and Photometry
The motion of a planet around its star and through the obscured regions of an observing instrument will affect the ability to measure a planet’s position and photometry. This depends to some extent on the orbit parameters of the planet and the observing system. If a planet goes inside the IWA or outside the OWA (outer-working angle) of the observing instrument, it will not be possible to get accurate astrometry or measure photometry. If the planet has an eccentric orbit, the cloud and surface properties may change substantially with orbital position. For example, varying instellation will cause different types of clouds to condense at different regions along the planet’s orbit (e.g., Refs. Marley et al. [43], Cahoy et al. [7]). It may also cause the surface to freeze over or thaw, therefore increasing or decreasing the surface albedo if the orbit takes a planet in/out of cooler and warmer conditions (e.g., Ref. Venkatesan et al. [69]).
Missed detections (e.g., when a planet falls inside/outside the working angles of the instrument) are also likely to increase confusion, because this may increase the number of orbits that can be fit to the points that are visible to the observer [53]. However, where a planet’s photometry can be measured, the likelihood ranking scheme would operate in the same manner. For the current model of photometric behavior and radiometric brightness, these keep-out zones and portions of the orbit where planets are not observable would not be input to the deconfuser. The user would only provide the observable points, and thus, the deconfuser would only fit orbits to the input points [53]. This would not change any of the outcomes of this work, as a user would have measured photometry for the input points that are provided to the deconfuser. Section 5 identifies future work to address cases where planets are unobservable in their orbit.
4.6 Comparing Photometry Ranking to Dynamical Assumptions
A significant fraction of known exoplanet systems exhibit eccentricities greater than 0.1. Of the systems with measured eccentricities, 46% have $e>0.1$ and 27% have $e>0.2$ [9]. Therefore, while dynamical stability is an important and necessary criterion for deconfusion, photometry provides a critical, independent discriminator for deconfusion decisions.
4.7 Observing Strategy
As was presented in Ref. Pogorelyuk et al. [53], the level of confusion depends on the number of observations and the system architecture. With three visits, the probability of confusion is still relatively high for high inclination systems ($40\%$ [53]). Increasing the number of visits to a system to four or five can significantly reduce ambiguity, but this comes at a cost to overall mission yield [62, 48]. Each observation of a system requires hours of integration time (dependent on the host star’s brightness and the required contrast) that use up mission resources and time that could be spent characterizing other systems. Due to these trade-offs, it is beneficial to overall mission yield to reduce the number of observations needed to characterize a system/planet and optimize the observing strategy for each system. Efficient deconfusion that leverages all available information from a series of observations is highly valuable in this regard.
4.8 Accuracy of Deconfuser Orbit Fits
To perform a thorough assessment of the deconfuser’s performance, we also calculated the accuracy of the semimajor axis fits returned by the deconfuser compared to the true semimajor axis for each planet, by comparing their percent difference. The accuracy of the fit for the semimajor axis is important for enabling follow-up observations and characterization of planets. The tolerance for orbital parameter accuracy defined by the NASA Standards Definition and Evaluation Team for Yield Analysis [47] is to better than ten percent accuracy. Of the planets in the low inclination systems simulated here, 64 percent of the top ranked orbit fits retrieved a semimajor axis within five percent of the true semimajor axis, 77 percent were retrieved to within ten percent of the true semimajor axis. All planets except one were fit to an accuracy of better than 25 percent. Of the medium inclination systems, 40 percent of the orbit fits retrieved the semimajor axis to within five percent of the true value, 70 percent to within ten percent of the true value, and 93 percent to within 25 percent of the true value. For the high inclination systems, the number of planets for which semimajor axes were retrieved to better than five percent of the true value dropped to 37 percent, 57 percent for constraints within ten percent, and 73 percent for constraints within 25 percent of the true semimajor axis. We emphasize that these percentages result after selection of the “best” orbit using the new photometry ranking scheme, where the best orbit is the one with the minimum percent difference between individual orbital parameters compared to the true system. In order to characterize potentially habitable planets with a future mission like HWO, observation schedulers will require accurate knowledge of whether a planet lies within the host star’s potential habitable zone and where it will be in the future in order to schedule the next observation.
4.9 Importance of Deconfusion for Observation Planning
Being able to deconfuse multiple orbit options using photometry will ultimately improve observation scheduling and enable characterization of exoplanet systems. It will also inform decisions about how to allocate mission time to studying planets of interest. For example, the confused orbit options for the inner two planets in the high inclination system ten give fits that change the number of planets with separations that fall within the boundaries of a Solar-type star’s liquid-water habitable zone (defined here with an inner edge at 0.836 AU and outer edge at 1.656 AU as in Ref. Kane and Gelino [30]). In the first option for the high inclination system ten, only the second planet in the system would fall within the liquid-water habitable zone. In the second option, both planet two and planet three would fall within the habitable zone. The difference between these orbital separations and architectures may inform how we determine follow-up observations on such a system. Observers may choose not to allocate valuable mission observing time and resources in a scenario where an orbit option places a planet outside of the habitable zone. The deconfuser with its photometry ranking scheme can support this decision making.
5 Summary & Future Work
In this work, we introduce a method and updated tool for using orbital phase and planetary brightness variation to deconfuse orbit combinations. We present a proof-of-concept analysis of a limited set of cases to explore the utility of planetary phase for planet-detection deconfusion. We apply the updated orbit ranking scheme to thirty representative highly confused systems containing three planets, categorized into three groups of ten across low, medium, and high inclination groupings. We adopt instrument parameters for the Roman Space Telescope Coronagraph Instrument to implement the photometric orbit-ranking scheme with realistic instrument values, but the results remain applicable to any direct imaging mission.
The analysis presented here shows that combining photometry with astrometry shows promise for improving deconfusion of multi-planet systems and enabling orbit differentiation. For the corner cases presented here, each system had multiple closely ranked orbit options using only relative astrometry, which results in confusion. The photometry ranking scheme helps to choose the system fit that most closely resembles the true simulated system in more than half of the representative cases across inclination groupings. Future work will explore the potential for improving the error rates on orbit ranking with photometry by setting thresholds for comparing system likelihoods and using more realistic planetary reflectance treatments.
On top of supporting observation scheduling and follow-up decisions by providing an estimate of where a planet resides in its system, orbit determination also supports placing better constraints on exoplanet properties. Ref. Salvador et al. [57] showed that a priori knowledge of a planet’s orbit supports accurate retrieval of the radius in reflected light observations. Orbit determination helps to break the degeneracy between planet radius and phase angle. Knowing a planet’s radius can help differentiate between clear and cloudy atmospheric scenarios, and supports placing constraints on atmospheric abundances of species like CH4 [8]. An unknown planet radius reduces the accuracy of the atmospheric retrieval for reflected light spectroscopy. Therefore, reducing the radius uncertainty on a planet through orbit determination is a useful prior for exoplanet characterization.
During direct imaging observations, planets may also be blocked by the inner or outer working angles of the observing instrument. This would result in missed detections when the planet falls within these obscured regions. Detections might also be missed if they are at a phase angle that is too dim to observe. In these cases, the probability of confusion for a system will be higher [53]. However, with photometric information, we may be able to reject or assign a lower ranking to geometrically valid orbits which would require the planet to be at a phase angle which is inconsistent with the detection’s brightness.
Future work will also account for the joint dependence of astrometry and photometry in the simulations. That is, the knowledge of a planet’s astrometry will be less accurate if its phase angle is higher and the planet is therefore dimmer, and vice versa. A detailed study of the effect of non-detections on confusion rates and the scheduling (and number) of visits per system would also directly inform yield calculations for future missions. Adjusting the observing cadence of a system may support reducing confusion that remains due to closely separated planet detections. Deviating from uniformly spaced observations, which were previously suggested for single-planet systems [26], might enable better distinction in separation and phase space. With additional orbit ranking metrics and simulations, we can provide a more complete assessment of confusion rates and their impact on yield for a mission like HWO.
Acknowledgments
S. N. H. would like to thank Eric Nielsen for helpful discussion on MCMC orbit-fitting and Vanessa Bailey for support in specifying the Roman Coronagraph Instrument parameters. The authors would like to thank those who reviewed the manuscript for their helpful feedback and suggestions. This research was carried out at the Massachusetts Institute of Technology and in part at the Jet Propulsion Laboratory, California Institute of Technology.
Appendix A Likelihood Ranking
Here we describe the likelihood ranking scheme developed for deconfusion with photometry. In this work, we consider each planet detection to be an independent event, as described in Sec. 2.2. Each planet detection will have a unique photon count that is dependent on factors including the orbital parameters of the system, the observation time, the background contribution, and the planet’s properties. The number of photons detected also depends on the parameters of the observing instrument, which motivates the inclusion of the detector model described in Secs. 2.1 and 2.2.
Here we describe the number of counts collected by the observing detector with a slightly simplified approach for clarity, adopting the observation time as $\Delta t$, the background counts as $d$, and the photon flux as $\phi$. The background here includes contributions from both the detector and astrophysical sources. The expected number of photons for a detection ($n_{p}$) is obtained by combining these possible contributions to the photon count, which can be described by Eq. (9).
$n_{p}(a,e,i,\Omega,\omega,M_{0})=\phi_{p}(a,e,i,\Omega,\omega,M_{0})\cdot\Delta t+d$ (9)
The orbital parameter definitions are listed in Table 4.